Suppose you have to distribute the numbers {1; 2; 3; ... ; 2n-1;2n} over n buckets. Show that there will always be at least one bucket with its sum of numbers to be >=2n+1.


1
Expert's answer
2019-10-29T14:37:35-0400

Let us assume that the sum of numbers in each bucket is less than 2n +1.

Where n is the number of buckets into which numbers ranging from 1 to 2n are distributed.

Let the sum of numbers in first bucket is s1.

Similarly there are s2,s3,s4,...,sn sums.

"s1+s2+...+sn = 1+2+...+(2n-1)+(2n)"

"s1 +s2+...+sn = (2\u00d7n)\u00d7(2\u00d7n+1)\/2"

"s1 + s2 +...+sn = n\u00d7(2\u00d7n+1)" ------(i)

According to our assumption

"s1<(2\u00d7n+1)"

"s2<(2\u00d7n+1)"

Similarly, "sn<(2\u00d7n+1)"

Adding up the above equations,we get ;

"s1 +s2+...+sn <(2\u00d7n+1)\u00d7n"

But this is a contradiction to the equation --(i)

So our assumption is wrong.

Thus,there will always be at least one bucket

with its sum of numbers to be >=2n+1.

(Proved)